OpenStudy (anonymous):

(Newton's Law of Motion) A 100 kg motorcycle carrying a 70 kg rider comes to a stop in 40 m from a speed of 50 km/h when its brakes are applied. Find the force exerted by the brakes and the force experienced by the rider.

OpenStudy (anonymous):

there are 2 ways of solving this question - 1-1/2mv^2 kinetic engery formula just subsitute the values 1/2 100+70x(50)^2 u get energy rite as u know wordone is in joules so is energy so take kinetic energy as workdone workdone =fxdistance f=wordone/distance 2 second method find acceleration v^2=u\[v ^{2}=u ^{2}+2as\] find acceleation f=MA find the force..

OpenStudy (anonymous):

use equation \[v ^{2}-u ^{2}=2as\] v=0m/s , u=50km/h= 125/9 m/sec s=40m you get accl =-25/144m/sec F=MA the force exerted by the brakes = Total mass\[\times\]accl=(100+70)-25/144 force experienced by the rider= 70\[\times\]-25/144

OpenStudy (anonymous):

v^2-u^2=2ad\[\rightarrow\]a=25.1m/s62 f=ma\[\rightarrow \f=170*25.1=4267 f=ma f=70*25.1=1757 n